OpenStudy (anonymous):

5u + 3v = -16 2u + v = -8 Does the negative sign change to positive or stay the same when I add -16 and -8?

OpenStudy (anonymous):

If so, -16 + -8 = 24 then 24/4 = 6 so v=6.

OpenStudy (anonymous):

5u + 3v = -16 (1) 2u + v = -8 (2) In order to solve the system, you need to multiply the second equation for 3 first

OpenStudy (anonymous):

I'm sorry, I don't get it...

OpenStudy (anonymous):

Just a moment I am trying to figure it out

OpenStudy (anonymous):

I think I know....

OpenStudy (anonymous):

In the previous one, it makes convenient that x terms are equal, In this case you don't have it ready!

OpenStudy (anonymous):

-6u - 3v = 24 correct?

OpenStudy (anonymous):

Wow, fast learner :)

OpenStudy (anonymous):

Now add the 2 equations to eliminate v, can you!

OpenStudy (anonymous):

the new system of equations are: 5v + 3v = -16 and -6 - 3v = 24 ... I got -8

OpenStudy (anonymous):

You're correct :) u = -8

OpenStudy (anonymous):

Just plug u= -8 back to the original equation

OpenStudy (anonymous):

I believe so... 5(-8) + 3v = -16 correct? -16/-40 ? that doesn't look right... lol

OpenStudy (anonymous):

Oh, use the original equation... 2(-8) + v = -8

OpenStudy (anonymous):

.5? 2 x -8 = -16 so -8/-16, right? This part always confuses me.

OpenStudy (anonymous):

2(-8) + v = -8 -16 + v = -8 --> v = -8 + 16 = 8

OpenStudy (anonymous):

I see, I keep trying to solve for v like I did for u but I have to add 16 to the other side like you did in the previous equation.

OpenStudy (anonymous):

Why don't you practice solve for V: 5u + 3v = -16

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

v=24

OpenStudy (anonymous):

5(-8) = -40, -16 +40 = 24 which is why I say v=24

OpenStudy (anonymous):

You need to practice writing in Algebra way: 5u + 3v = -16 5 (-8) + 3v = -16

OpenStudy (anonymous):

-> 3v = -16 + 40 = 24 => v = 24/3 = 8

OpenStudy (anonymous):

You're right, I do thank you...

OpenStudy (anonymous):

I noticed that, it was a little sloppy on my paper as well.

OpenStudy (anonymous):

Ahh, that makes sense! Okay. Thank you!

OpenStudy (anonymous):

Feel overwhelmed now ;)

OpenStudy (anonymous):

No, you've helped a lot thank you.

OpenStudy (anonymous):

Keep practice, you learn fast :)

OpenStudy (anonymous):

I will, thank you so much. :)